A car moves a distance of 200 m. It covers the first half of the distance at speed 40 km/h and the second half of the distance at speed v. If the average speed is 48 km/h, then the value of v is
 

Grade:9

1 Answers

Arun
4 years ago
48=   200/((100/40)+(100/v) )
    or 1/40+1/v=  2/48=1/24
    or  1/v=1/24-1/40=2/120=1/60
    or v=60km/h
    [Alt:  v_av  = (2v_1 v_2)/(v_1+v_2 ) => 48= (2×40×v)/(40+v)   =>  v=60 km/h]